Predict the sign of the entropy change, Δ S°, for each of the reaction displayed. Drag the appropriate items to their respective bins.

\(\mathrm{Ag}^{+}(\mathrm{aq})+\mathrm{Br}^{-}(\mathrm{aq}) \rightarrow \mathrm{AgBr}(\mathrm{s})\)

\(\mathrm{Ca}(\mathrm{OH})_{2}(\mathrm{~s}) \rightarrow \mathrm{CaO}(\mathrm{s})+\mathrm{H}_{2} \mathrm{O}(\mathrm{g})\)

\(2 \mathrm{~N}_{2} \mathrm{O}(\mathrm{g}) \rightarrow 2 \mathrm{~N}_{2}(\mathrm{~g})+\mathrm{O}_{2}(\mathrm{~g})\)

\(2 \mathrm{Na}(\mathrm{s})+\mathrm{Cl}_{2}(\mathrm{~g}) \rightarrow 2 \mathrm{NaCl}(\mathrm{s})\)

\(\mathrm{C}_{4} \mathrm{H}_{8}(\mathrm{~g})+6 \mathrm{O}_{2}(\mathrm{~g}) \rightarrow 4 \mathrm{CO}_{2}(\mathrm{~g})+4 \mathrm{H}_{2} \mathrm{O}(\mathrm{g})\)

\(\mathrm{I}_{2}(\mathrm{~s}) \rightarrow \mathrm{I}_{2}(\mathrm{~g})\)


Part A Predict the sign of the entropy change, ΔS° , for each of the reaction displayed. Drag the appropriate items to their respective bins. Ag+(aq) + Br-(aq) → AgBr(s) | Ca(OH)2(s) → CaO(s) + H2O(g) 2N20(g) → 2N2(g) + O2(g) | | 2Na(s) + Cl2(g) → 2NaCl(s) C,Hs(g) + 602(g) 4CO2(g) + 4H2O(g) | 12(s) → 12(g) Positive Negative reset help reset?help
Answer #1
Concepts and reason

The change of entropy (ΔSo)\left( {\Delta {{\rm{S}}^{\rm{o}}}} \right) is the difference in entropy of the products and the entropy of reactants.

ΔSo=ΔSo(products)ΔSo(reactants)\Delta {{\rm{S}}^{\rm{o}}} = \Delta {{\rm{S}}^{\rm{o}}}\left( {{\rm{products}}} \right) - \Delta {{\rm{S}}^{\rm{o}}}\left( {{\rm{reactants}}} \right)

Its sign is positive if the number of gaseous particles on the product side is greater than the number of gaseous particles on the reactant side.

Fundamentals

The entropy is the amount of disorderness. In solid state, the particles are closely packed; hence, the disorderness is negligible. Liquids have more disorderness when compared to solids, but it is negligible when compared to gaseous particles.

Consider the following reaction:

Ag+(aq)+Br(aq)AgBr(s)\,{\rm{A}}{{\rm{g}}^ + }\left( {aq} \right) + {\rm{B}}{{\rm{r}}^ - }\left( {aq} \right) \to {\rm{AgBr}}\left( s \right)

Here, products have less entropy than reactants.

Hence, for this reaction, the sign ΔSo\Delta {{\rm{S}}^{\rm{o}}} is negative.

Consider the following reaction:

Ca(OH)2(s)CaO(s)+H2O(g){\rm{Ca}}{\left( {{\rm{OH}}} \right)_2}\left( s \right) \to {\rm{CaO}}\left( s \right) + \,{{\rm{H}}_{\rm{2}}}{\rm{O}}\left( g \right)

Here, products have more entropy than reactants.

Hence, for this reaction, the sign ΔSo\Delta {{\rm{S}}^{\rm{o}}} is positive.

Consider the following reaction:

2N2O(g)2N2(g)+O2(g){\rm{2}}{{\rm{N}}_{\rm{2}}}{\rm{O}}\left( g \right) \to 2{{\rm{N}}_2}\left( g \right) + \,{{\rm{O}}_2}\left( g \right)

Here, products have more entropy than reactants.

Hence, for this reaction, the sign ΔSo\Delta {{\rm{S}}^{\rm{o}}} is positive.

Consider the following reaction:

2Na(s)+Cl2(g)2NaCl(s)\,2{\rm{Na}}\left( s \right) + \,{\rm{C}}{{\rm{l}}_2}\left( g \right) \to 2{\rm{NaCl}}\left( s \right)

Here, products have less entropy than reactants.

Hence, for this reaction, the sign ΔSo\Delta {{\rm{S}}^{\rm{o}}} is negative.

Consider the following reaction:

C4H8(g)+6O2(g)4CO2(g)+4H2O(g)\,{{\rm{C}}_4}{{\rm{H}}_8}\left( g \right) + 6{\rm{ }}{{\rm{O}}_2}\left( g \right) \to \,4{\rm{C}}{{\rm{O}}_2}\left( g \right){\rm{ + 4}}{{\rm{H}}_{\rm{2}}}{\rm{O}}\left( g \right)\,\,\,

Here, products have more entropy than reactants.

Hence, for this reaction, the sign ΔSo\Delta {{\rm{S}}^{\rm{o}}} is positive.

Consider the following reaction:

I2(s)I2(g){{\rm{I}}_2}\left( s \right) \to {{\rm{I}}_2}\left( g \right)

Here, products have more entropy than reactants.

Hence, for this reaction, the sign ΔSo\Delta {{\rm{S}}^{\rm{o}}} is positive.

Ans:

Drag the appropriate items to their boxes as follows:

Positive
Negative
Agt (aq)+ Br (aq) -AgBr(s)
Ca(OH)2(s)Ca(s) + H2O(g)
2N2O(g) 2N2(g)O2(g)
2Na(s)+Cl2(g) - 2NAC1(s)
C4H&(g)+60